''After five years age of mother will be 3 times the age of her son''. Represent above statement as a linear equation in two variables.represent in form of linear equation

Answer by stanbon(75887)

''After five years age of mother will be 3 times the age of her son''. Represent above statement as a linear equation in two variables.represent in form of linear equation
-----------------
m+5 = 3(s+5)
m+5 = 3s + 15
m-3s = 10
---------------
Cheers,
Stan H.
------------